OpenStudy (kitkat1):

The work of a student to solve a set of equations is shown: Equation A: y = 15 – 2z Equation B: 2y = 3 – 4z Step 1: –2(y) = –2(15 – 2z) [Equation A is multiplied by –2.] 2y = 3 – 4z [Equation B] Step 2: –2y = 15 – 2z [Equation A in Step 1 is simplified.] 2y = 3 – 4z [Equation B] Step 3: 0 = 18 – 6z [Equations in Step 2 are added.] Step 4: 6z = 18 Step 5: z = 3 In which step did the student first make an error? Step 2 Step1 Step 4 Step 3

OpenStudy (kitkat1):

I will give medals!

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Recall the distributive property of multiplication: a(b+c)=ab+ac.

OpenStudy (kitkat1):

Im pretty sure its step 2

OpenStudy (kitkat1):

Because there he used subtraction instead of division I need to confirm that

OpenStudy (kitkat1):

Someone Please help!

OpenStudy (skullpatrol):

y = 15 – 2z –2y = –2•( 15 – 2z )

OpenStudy (kitkat1):

So im i right with the step 2?

OpenStudy (skullpatrol):

You're right! It is step 2 :-)

OpenStudy (kitkat1):

okay thank you can u help me with one more?

OpenStudy (kitkat1):

Variable c is 5 more than variable a. Variable c is also 3 less than variable a. Which pair of equations best models the relationship between c and a? a = c + 5 a = 3c –3 c = a – 5 c = a + 3 a = c – 5 a = 3c + 3 c = a + 5 c = a – 3

OpenStudy (kitkat1):

Im pretty sure its the first one

OpenStudy (kitkat1):

can u confirm?

OpenStudy (skullpatrol):

The -2 was NOT distributed correctly.

OpenStudy (kitkat1):

okay can u confirm the 2nd one please?

OpenStudy (skullpatrol):

The question asks for which pair.

OpenStudy (skullpatrol):

That means two equations, right?

OpenStudy (skullpatrol):

Variable c is 5 more than variable a: c = 5 + a

OpenStudy (skullpatrol):

Variable c is 3 less, than variable a: c = a – 3

OpenStudy (skullpatrol):

You should email your teacher and tell them that there is no relationship between a and c because the equations have no solution.

OpenStudy (skullpatrol):

5+a=a-3 gives 5=-3 which is not true